Consider the non-linear inverse demand function, p= -q2-q+I. Given the total cost function: TC =9q2+2q+
i. Find the marginal cost (MC) and average cost (AC) as functions of q
ii. Find the output at which average cost is minimized
1
Expert's answer
2020-04-30T10:03:29-0400

i. Find the marginal cost (MC) and average cost (AC) as functions of q

The total cost function is:



"TC =9q^2+2q+ 81"

The marginal cost is:



"MC = \dfrac{TC}{\Delta Q}= 18q + 2"

The average cost is:



"AC = \dfrac{TC}{q} = 9q + 2 + \dfrac{81}{q}"


ii. Find the output at which average cost is minimized


The marginal cost crosses the average cost at its minimum. Thus, at the minimum of the average cost, "MC = AC."



"18q + 2 = 9q + 2 + \dfrac{81}{q}"

"9q = \dfrac{81}{q}"

"q^2 = 9"

"q^* = \sqrt{9} = 3"